Find the future value in two years of $100 that is deposited in an account, which pays 12%, compounded monthly.

$118.80

$112.70

$125.40

$126.97

$160.00
Answer #1

Amount = P(1 + r/n)nt

Here,

n = No. of times interest compounded

Amount = 100 (1 + 0.12/12)12*2

= 100 (1.01)24

Amount = $126.97